the nut shoppe has 10 kg of mixed cashews and pecans,which sell for $8.90 per kilogram.cashews alone sell for $8 per kilogram,and pecans sell for $9 per kilogram.how many kilograms of each are in the mix?
Answer by checkley71(8403)

8X+9(10-X)=10*8.9
8X+90-9X=89
-X=89-90
-X=-1
X=1 KG OF CASHEWS.
10-1=9 KG OF PECANS.
PROOF
8*1+9*9=89
8+81=89
89=89